# Towards Robust Object Detection: Identifying and Removing Backdoors via Module Inconsistency Analysis

### v1

@@ Abstract.
Object detection models have been widely adopted in various security-critical applications, such as autonomous driving and video surveillance. However, the complex architectures of these models also make them vulnerable to backdoor attacks, where maliciously trained models behave normally on clean inputs but produce targeted misclassifications when triggered by specific patterns. Existing backdoor defense techniques, primarily designed for simpler models like image classifiers, often fail to effectively detect and remove backdoors in object detectors while preserving model performance. In this work, we propose a novel backdoor defense framework tailored to the unique characteristics of object detection models (Liang et al., 2024b). Our key observation is that a backdoor attack often causes significant inconsistencies between the behaviors of local modules, such as the Region Proposal Network (RPN) and the classification head. By quantifying and analyzing these inconsistencies, we develop an effective algorithm to detect the presence of backdoors. Furthermore, we find that the inconsistent module is usually the main source of the backdoor behavior. Exploiting this insight, we propose a simple yet effective backdoor removal method, which localizes the affected module, resets its parameters, and fine-tunes the model on a small set of clean data. Extensive experiments with multiple state-of-the-art object detectors demonstrate that our method can successfully detect and remove backdoors, achieving an improvement of 90% in the backdoor removal rate over the fine-tuning baseline while limiting the accuracy loss on clean data to less than 4%. To the best of our knowledge, this work represents the first effort to develop a dedicated backdoor defense framework for object detection models, addressing the unique challenges and limitations of existing techniques in this context. Our work sheds new light on the unique challenges and opportunities in defending object detection models against backdoor attacks.
@@ 1. Introduction
Object detection plays a pivotal role in a wide range of multimedia applications, such as video surveillance (Raghunandan et al., 2018), autonomous driving (Feng et al., 2020), and face recognition (Hu et al., 2015). However, the increasing deployment of deep learning-based object detectors in real-world scenarios has raised growing concerns about their security vulnerabilities, especially the emerging threats of backdoor attacks. Backdoor attacks refer to the malicious manipulation of a model’s behavior by injecting hidden triggers during the training process, which can cause severe consequences like unauthorized access and privacy leakage (et al., 2022a, 2023j, 2023b, 2023a), thereby jeopardizing the trustworthiness and reliability of multimedia systems heavily relying on object detection.
@@
Despite the severity of backdoor threats in object detection models, existing research efforts primarily focus on developing novel attack strategies (Ma et al., 2022a; et al., 2024c),  (Ma et al., 2022b), while the defense aspect remains largely underexplored. To the best of our knowledge, only a few recent works (Cheng et al., 2024),  (Shen et al., 2024) have made initial attempts to detect backdoors in object detectors, leveraging techniques like activation clustering and gradient analysis. However, these methods require a large number of clean samples or rely on certain assumptions about the backdoor patterns, which may not always hold in practice. Moreover, effective techniques for backdoor removal in object detection models are still missing in the literature, leaving a significant gap in the defense pipeline.
@@ 2.1. Backdoor Attacks
The seminal work on backdoor attacks (Liang et al., 2023; et al., 2023i, e, 2024b, 2024d) in image classification is BadNets (Gu et al., 2019), which injects a backdoor into a model by poisoning a portion of the training data with a specific trigger pattern (e.g., a black square) and the target label. The model trained on this poisoned dataset will behave normally on clean inputs, but misclassify any input containing the trigger to the target label. The proposal of BadNets has sparked extensive research on backdoor attacks in the academic community, leading to a diverse array of attack techniques. Blended (Chen et al., 2017) replaces the prominent black square trigger with a more stealthy transparent cartoon sticker. SSBA (Li et al., 2021a) takes inspiration from the classical image steganography (Cheddad et al., 2010) technique and embeds the trigger into the image using steganographic methods, making the trigger imperceptible to human eyes. AdvDoor (Zhang et al., 2021), on the other hand, borrows the idea from adversarial attacks (et al., 2021, 2020, 2018, 2022c, 2022b, 2023n, 2023k, 2023c, 2023h, 2023d, 2023l, 2024a) and uses adversarial perturbations as a backdoor trigger. FaceHack (Sarkar et al., 2021) targets face recognition systems by using facial paintings as triggers. Wanet (Nguyen and Tran, 2021) employs subtle warping of the edge of the object as a trigger, while low frequency (Zeng et al., 2021) inserts the trigger in the frequency domain.
As the research on backdoor attacks against image classification models becomes more mature (Wu et al., 2023a), researchers have started to turn their attention to the vulnerability of object detection models. Chan (Chan et al., 2022) are the first to propose four types of backdoor attacks specifically designed for object detection models, this work reveals that object detection models are equally threatened by backdoor attacks. Luo (Luo et al., 2023) further investigates the object disappearance attack. They conduct a detailed study on this specific attack and demonstrate that even using the simplest attack method, some basic defense (et al., 2023m, g, f) techniques such as fine-tuning and fine-pruning are still ineffective against this type of attack. This work highlights the severity of the backdoor threat in object detection models and the inadequacy of existing defense methods. Taking into account real-world scenarios, Ma  (Ma et al., 2023) proposes a clean label backdoor attack method called TransCAB, which uses natural triggers. TransCAB employs a Transformer to model the relationship between object instances and object appearances in natural images, generating realistic poisoned data containing triggers to compromise the model.
@@ 2.2. Backdoor Defenses
In the image classification domain, a variety of backdoor defense techniques (et al., 2022d; Liang et al., 2024a) have been proposed, showing promising results in detecting and mitigating backdoor attacks. Fine-Pruning (FP) (Liu et al., 2018) is one of the early methods that combines network pruning and fine-tuning to remove backdoor-related neurons while preserving the model’s accuracy on clean data. Later, Neural Attention Distillation (NAD) (Li et al., 2021c) proposes to first fine-tune a clean teacher model using clean samples and then transfer the attention maps of the teacher model to the backdoored student model to erase the backdoor effect. Building upon NAD, FTT (Zhang et al., 2024) further improves the model’s accuracy recovery and redesigns the defense against advanced attacks. However, it is worth noting that most of these defense techniques are primarily evaluated on traditional image classification models. Their applicability and effectiveness on object detection models, which have significantly different architectures and learning objectives, remain to be further investigated.

### v2

@@ Abstract.
Object detection models have been widely adopted in various security-critical applications, such as autonomous driving and video surveillance. However, the complex architectures of these models also make them vulnerable to backdoor attacks, where maliciously trained models behave normally on clean inputs but produce targeted misclassifications when triggered by specific patterns. Existing backdoor defense techniques, primarily designed for simpler models like image classifiers, often fail to effectively detect and remove backdoors in object detectors while preserving model performance. In this work, we propose a novel backdoor defense framework tailored to the unique characteristics of object detection models (Liang et al., 2024d). Our key observation is that a backdoor attack often causes significant inconsistencies between the behaviors of local modules, such as the Region Proposal Network (RPN) and the classification head. By quantifying and analyzing these inconsistencies, we develop an effective algorithm to detect the presence of backdoors. Furthermore, we find that the inconsistent module is usually the main source of the backdoor behavior. Exploiting this insight, we propose a simple yet effective backdoor removal method, which localizes the affected module, resets its parameters, and fine-tunes the model on a small set of clean data. Extensive experiments with multiple state-of-the-art object detectors demonstrate that our method can successfully detect and remove backdoors, achieving an improvement of 90% in the backdoor removal rate over the fine-tuning baseline while limiting the accuracy loss on clean data to less than 4%. To the best of our knowledge, this work represents the first effort to develop a dedicated backdoor defense framework for object detection models, addressing the unique challenges and limitations of existing techniques in this context. Our work sheds new light on the unique challenges and opportunities in defending object detection models against backdoor attacks.
@@ 1. Introduction
Object detection plays a pivotal role in a wide range of multimedia applications, such as video surveillance (Raghunandan et al., 2018), autonomous driving (Feng et al., 2020), and face recognition (Hu et al., 2015). However, the increasing deployment of deep learning-based object detectors in real-world scenarios has raised growing concerns about their security vulnerabilities, especially the emerging threats of backdoor attacks. Backdoor attacks refer to the malicious manipulation of a model’s behavior by injecting hidden triggers during the training process, which can cause severe consequences like unauthorized access and privacy leakage (Liang et al., 2022b; enhancing face obfuscation guided by semantic-aware attribution maps, 2023; Guo et al., 2023; Dong et al., 2023), thereby jeopardizing the trustworthiness and reliability of multimedia systems heavily relying on object detection.
@@
Despite the severity of backdoor threats in object detection models, existing research efforts primarily focus on developing novel attack strategies (Ma et al., 2022a; Li et al., 2024),  (Ma et al., 2022b), while the defense aspect remains largely underexplored. To the best of our knowledge, only a few recent works (Cheng et al., 2024),  (Shen et al., 2024) have made initial attempts to detect backdoors in object detectors, leveraging techniques like activation clustering and gradient analysis. However, these methods require a large number of clean samples or rely on certain assumptions about the backdoor patterns, which may not always hold in practice. Moreover, effective techniques for backdoor removal in object detection models are still missing in the literature, leaving a significant gap in the defense pipeline.
@@ 2.1. Backdoor Attacks
The seminal work on backdoor attacks (Liang et al., 2023b; Liu et al., 2023d, b; Liang et al., 2024a, b) in image classification is BadNets (Gu et al., 2019), which injects a backdoor into a model by poisoning a portion of the training data with a specific trigger pattern (e.g., a black square) and the target label. The model trained on this poisoned dataset will behave normally on clean inputs, but misclassify any input containing the trigger to the target label. The proposal of BadNets has sparked extensive research on backdoor attacks in the academic community, leading to a diverse array of attack techniques. Blended (Chen et al., 2017) replaces the prominent black square trigger with a more stealthy transparent cartoon sticker. SSBA (Li et al., 2021a) takes inspiration from the classical image steganography (Cheddad et al., 2010) technique and embeds the trigger into the image using steganographic methods, making the trigger imperceptible to human eyes. AdvDoor (Zhang et al., 2021), on the other hand, borrows the idea from adversarial attacks (Liang et al., 2021, 2020; Wei et al., 2018; Liang et al., 2022c, a; Wang et al., 2023; Liu et al., 2023a; He et al., 2023b; Liu et al., 2023e; He et al., 2023a; Muxue et al., [n. d.]; Lou et al., 2024) and uses adversarial perturbations as a backdoor trigger. FaceHack (Sarkar et al., 2021) targets face recognition systems by using facial paintings as triggers. Wanet (Nguyen and Tran, 2021) employs subtle warping of the edge of the object as a trigger, while low frequency (Zeng et al., 2021) inserts the trigger in the frequency domain.
As the research on backdoor attacks against image classification models becomes more mature (Wu et al., 2023a), researchers have started to turn their attention to the vulnerability of object detection models. Chan (Chan et al., 2022) are the first to propose four types of backdoor attacks specifically designed for object detection models, this work reveals that object detection models are equally threatened by backdoor attacks. Luo (Luo et al., 2023) further investigates the object disappearance attack. They conduct a detailed study on this specific attack and demonstrate that even using the simplest attack method, some basic defense (Sun et al., 2023; Liu et al., 2023c; Liang et al., 2023a) techniques such as fine-tuning and fine-pruning are still ineffective against this type of attack. This work highlights the severity of the backdoor threat in object detection models and the inadequacy of existing defense methods. Taking into account real-world scenarios, Ma  (Ma et al., 2023) proposes a clean label backdoor attack method called TransCAB, which uses natural triggers. TransCAB employs a Transformer to model the relationship between object instances and object appearances in natural images, generating realistic poisoned data containing triggers to compromise the model.
@@ 2.2. Backdoor Defenses
In the image classification domain, a variety of backdoor defense techniques (Wang et al., 2022; Liang et al., 2024c) have been proposed, showing promising results in detecting and mitigating backdoor attacks. Fine-Pruning (FP) (Liu et al., 2018) is one of the early methods that combines network pruning and fine-tuning to remove backdoor-related neurons while preserving the model’s accuracy on clean data. Later, Neural Attention Distillation (NAD) (Li et al., 2021c) proposes to first fine-tune a clean teacher model using clean samples and then transfer the attention maps of the teacher model to the backdoored student model to erase the backdoor effect. Building upon NAD, FTT (Zhang et al., 2024) further improves the model’s accuracy recovery and redesigns the defense against advanced attacks. However, it is worth noting that most of these defense techniques are primarily evaluated on traditional image classification models. Their applicability and effectiveness on object detection models, which have significantly different architectures and learning objectives, remain to be further investigated.